# Structural fluctuations in active glasses

### v2

@@ I Introduction
When the active matter is in sufficiently dense states, it often exhibits glassy dynamics [6, 7, 8]. This phenomenon has received considerable attention in biological physics, especially in studies of the collective motion of confluent epithelial cells that regulate wound healing, cancer metastasis, and other processes [6, 9, 10, 11, 12, 13, 14, 15], the diffusion and viscoelasticity of proteins in bacterial cytoplasm [16, 17], and the dynamics and rheology of dense bacterial suspensions [18, 19]. Correspondingly, many numerical and theoretical models of dense active matter have been proposed and studied [20, 21, 22]. Among them, the simplest is presumably the self-propelled particles. Previous works have established that this model shows canonical glassy dynamics when the density is high enough or the active force is weak enough, where the abrupt slowdown of the microscopic dynamics takes place, the dynamic heterogeneity emerges, and the system eventually vitrifies into a glass state called the active glass. However, several non-canonical dynamical behaviors have also been observed, such as a change in the glass transition density, violation of the fluctuation-dissipation theorem, and the emergence of novel collective dynamics [20, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Extensions of equilibrium theories of the glass transition have been proposed to understand these behaviors [32, 27, 30, 26, 33], but they are still the subject of intense debate.
@@
As mentioned above, previous studies have focused on the properties of dense active matter in their liquid state, while little is known about the properties of active glasses in their solid state. The present study aims to shed light on the structural fluctuations in active glasses. In passive glasses, particles are driven solely by thermal agitations, and their structural fluctuations are determined by their vibrational normal modes. Glasses have a large number of non-phonon normal modes, which differ significantly from crystalline solids where the normal modes are longitudinal and transverse phonons, and their vibrational density of states (vDOS) follows the Debye law $g(\omega)\propto\omega^{d-1}$ (where $d$ is the spatial dimension). As a result, there exist excess vibrational states over the Debye law, which is observed by a large peak in the reduced vDOS $g(\omega)/\omega^{d-1}$, referred to as the boson peak [34, 35, 36]. Additionally, at much lower frequencies, the vibrational normal modes become a mixture of phonon vibrations and quasi-localized vibrations [37, 38, 39, 40]. Within the harmonic description, structural fluctuations in passive glasses can be described in terms of the thermal excitations of these characteristic normal modes, all of which follow the equipartition law. Even in anharmonic dynamics at finite temperatures, the quasi-localized vibrations can be used to indicate structural rearrangements [41, 42]. Therefore, to understand the structural fluctuations in active glasses, it is essential to examine how the active forces excite the characteristic normal modes in glasses.
@@
In the present work, we study active glasses by employing three different models: Active Ornstein Uhlenbeck Particles (AOUPs) [43], Active Brownian Particles (ABPs), and Run and Tumble Particles (RTPs) [44]. To characterize structural fluctuations, we consider four observables: average excitation energy of the normal modes, mean-squared displacements (MSD), structure factor of longitudinal and transverse displacement fields, and non-Gaussian parameters (NGP). Firstly, we demonstrate that the former three observables can be described in terms of the two-time correlations of active forces and the normal modes of glasses. Next, we formulate the two-time correlations of the active forces to obtain general formulas of the three observables. Additionally, we prove that AOUPs, ABPs, and RTPs have the same two-time correlation when the active forces are properly reparameterized. We then apply the obtained formulas to calculate the observables for a typical configuration of active glasses. Our central observation is that active forces behave like random noise for low-frequency normal modes, while they behave like constant forces for high-frequency normal modes. This results in thermal excitations of low-frequency modes on one hand and quasi-static dynamics of high-frequency modes on the other, both of which determine nature of the structural fluctuations in active glasses. Finally, we study the NGP and find that fluctuations are mostly Gaussian, except for normal modes at the high-frequency edge. It should be noted that a previous study conducted by Henke et al. [45] investigated the velocity fields in confluent cells, both experimentally and theoretically. In their study, they modeled the cells using the ABPs, and demonstrated that the excitations of the normal modes in the model were able to accurately describe the velocity fields of cells as measured by experiments. The present work provides a more thorough analysis of the displacement fields, which complements and generalizes the work done by Henke et al. [45].
@@ II.2 Equation of motion
Because we assumed the mechanical stability of the configuration ${\bf\it r}_{0}$, the dynamical matrix $\mathcal{M}$ is positive semi-definite and has two translational zero modes. By ignoring these zero modes, we can avoid divergences in the displacement ${\bf\it u}$ over a long time and it only fluctuates around ${\bf\it u}={\bf\it 0}$. In this study, we solve Eq. (5) to investigate the fluctuation dynamics in the steady state. We consider three different active forces, each corresponding to a different active glass model.
@@ II.2.1 Active Ornstein Uhlenbeck Particles
In the AOUPs, the active force follows
@@ II.2.2 Active Brownian Particles
In the ABPs, the particle tries to maintain a constant speed while their direction changes constantly. The active force acting on ABPs is described by
@@ II.2.3 Run and Tumble Particles
In the RTPs, the particle attempts to move at a constant speed, similar to the ABPs. However, their directions change abruptly and randomly. The force vector, ${\bf\it f}$, is described by Eq. (8). Meanwhile, each angle $\theta_{i}$ changes randomly in the range of $0\leq\theta_{i}<2\pi$ with a probability per unit time $P_{\text{\rm tumble}}=1-e^{-R_{\rm tumble}}$, where $R_{\rm tumble}$ is a constant rate. Hereafter, the active glass model that uses this active force is referred to as the RTP glass.
@@ II.4 Observables
Here, $\vec{u}_{\vec{k}}$ is the Fourier transform of the displacement field ${\bf\it u}=(\vec{u}_{1},\vec{u}_{2},\cdots,\vec{u}_{N})$,
$\vec{u}_{\vec{k}}=\sum_{j=1}^{N}\vec{u}_{j}\exp(-i\vec{k}\cdot\vec{r}_{j}),$ (18)
and the longitudinal and transverse parts are defined as $\vec{u}_{L,\vec{k}}=\left(\hat{\vec{k}}\cdot\vec{u}_{\vec{k}}\right)\hat{\vec{k}}$ and $\vec{u}_{T,\vec{k}}=\vec{u}-\vec{u}_{L,\vec{k}}$, where $\hat{\vec{k}}=\vec{k}/k$ is the unit vector in the direction of $\vec{k}$, and $k=|\vec{k}|$. Using Eq. (12), we can express these correlation functions in terms of the eigenmodes as
@@
×$\times\left(\hat{\vec{k}}\cdot\vec{e}_{l,i}\right)\left(\hat{\vec{k}}\cdot\vec{e}_{m,j}\right)\exp\left[i\left(\vec{k}\cdot\left(\vec{r}_{0,j}-\vec{r}_{0,i}\right)\right)\right],$
@@
×$\times\left[\vec{e}_{l,i}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}_{l,i}\right)\right]\left[\vec{e}_{m,j}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}_{m,j}\right)\right]$
×$\times\exp\left[i\left(\vec{k}\cdot\left(\vec{r}_{0,j}-\vec{r}_{0,i}\right)\right)\right],$
where $\vec{r}_{0,i}$ denotes the position of the particle $i$ in the inherent structure ${\bf\it r}_{0}$.
@@
Finally, we quantify the non-Gaussianity of the fluctuation dynamics by two NPGs. One NGP is defined for the active force as
@@ III.2 ABP glass
The time correlations on the right-hand side of Eq. (26) can be evaluated as follows. Since ${\bf\it\theta}(t)$ follows the Wiener process Eq.(9), the conditional probability $P({\bf\it\theta},t|{\bf\it\theta}^{\prime},t^{\prime})$, which expresses the probability that the angle vector is ${\bf\it\theta}$ at time $t$ given that it is ${\bf\it\theta}^{\prime}$ at time $t^{\prime}$, is expressed as
@@
$=\gamma^{2}v_{0}^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\{e_{l,i}^{x}e_{m,j}^{x}+e_{l,i}^{y}e_{m,j}^{y}\}\frac{1}{2}e^{-D|t-t^{\prime}|}\delta_{ij},$
@@ III.3 RTP glass
The active force $f_{l}$ in the RTP glass shares the same expression Eq. (25) with the ABP glass. Therefore, the time correlations can be written as in Eq. (26). A difference between the RTPs and ABPs comes from the time evolution of ${\bf\it\theta}(t)$. In the RTP glass, ${\bf\it\theta}(t)$ follows the Poisson process, and the conditional probability is given by [47]
@@ III.4 Comparison of three active forces
Moving forward, we will utilize the parameters $(D,\tau)$ to identify the three active forces. Specifically, $(v_{0}=\sqrt{2D/\tau\gamma},D_{R}=1/\tau)$ is used to describe the ABP glass, while $(v_{0}=\sqrt{2D/\tau\gamma},R_{\rm tumble}=1/\tau)$ is used to describe the RTP glass. This parameterization will enable us to compare the three active glasses in a fair manner using $(D,\tau)$.
@@ IV Numerical model
We conduct numerical simulations by solving the linearized equation of motion, Eq. (5), with three different active forces, all of which begin from the same mechanically stable configuration ${\bf\it r}_{0}$. We adopted the Predictor-Corrector Method for the numerical integration. The active forces were computed using different methods depending on the type of system. For the AOUP glass, we use the Gillespie method [50], while for the ABP glass, we use the Euler-Maruyama method. For the RTP glass, we select the update event of the active forces to occur with a probability $P_{\text{tumble}}=1-e^{-R_{\text{tumble}}\Delta t}$ at each time step $\Delta t$, and then the direction of the active forces was chosen according to the uniform distribution. In all three cases, we perform simulations for a sufficiently long period to achieve stationary states before running the production runs.
@@ V.1 Time correlation function of $c_{l}$
This outcome is identical to the one for thermal glasses, where $D$ denotes temperature and $\gamma\omega_{l}^{2}$ symbolizes the stiffness of the mode $l$. It demonstrates that under the condition $\tau\ll\omega_{l}^{-2}$ which we term the “thermal condition” from now on, the active forces function like a random white noise for the mode $l$, and the dynamics along it are equivalent to those in thermal systems.
@@
It is important to note that the typical amplitude of $c_{l}$ is heavily influenced by these conditions, as discussed in detail later In the thermal condition, we observe that $\langle c_{l}^{2}\rangle\propto\omega_{l}^{-2}$, which is consistent with the equipartition law. However, in the quasi-static condition, $\langle c_{l}^{2}\rangle\propto\omega_{l}^{-4}$, meaning that displacements along the high-frequency modes are significantly suppressed.
@@ V.4 Spatial correlation of displacements
We now discuss the spatial correlations of the longitudinal and transverse displacement fields, denoted as $S_{L}(k)$ and $S_{T}(k)$, respectively. These correlation functions are calculated in our numerical simulations using Eq. (17). In Figs. 4(a) and 5(a), we present $S_{L}(k)$ and $S_{T}(k)$ for three different active glasses with varying persistent times $\tau$. When $\tau=10^{-8}$, $S_{L}(k)$ and $S_{T}(k)$ exhibit a plateau in a wide range of $k$. This is consistent with the known results for thermal glasses [46]. However, as $\tau$ is increased, both $S_{L}(k)$ and $S_{T}(k)$ become smaller at large $k$, indicating that short-wavelength excitations are strongly suppressed. This indicates that the fluctuations of active glasses exhibit apparent collective motions as the long-wavelength excitations are emphasized. These apparent collective motions only appear when the persistence time $\tau$ is sufficiently large. Notably, the above observations for all three active glasses are identical.
@@
×$\times\left(\hat{\vec{k}}\cdot\vec{e_{l,i}}\right)\left(\hat{\vec{k}}\cdot\vec{e_{l,j}}\right)\exp\left[i\left(\vec{k}\cdot\left(\vec{r}_{0,j}-\vec{r}_{0,i}\right)\right)\right],$
@@
×$\times\left[\vec{e}_{l,i}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}_{l,i}\right)\right]\left[\vec{e}_{l,j}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}_{l,j}\right)\right]$
@@
Here, we focus on the low-wavenumber regime in Eq. (50). We can use conventional elasticity theory to approximate the eigenmodes in this regime. Specifically, in the two-dimensional continuum solid, we can approximate the dynamical matrix Eq. (3) as the operator ${\frac{1}{\gamma\rho}\left(K\nabla\nabla+G\nabla^{2}\right)}$, which acts on the continuum displacement fields. Here, $K$ and $G$ are the bulk and shear modulus, respectively, and $\rho=N/V$ is the number density. We can solve the eigenvalue problem of this operator and obtain the eigenmodes specified by $l=\vec{k}\alpha$ where $\vec{k}$ is the wavevector and $\alpha$ is the polarization with $\alpha=L$ and $T$ denoting the longitudinal and the transverse modes, respectively. The eigenfrequency can be expressed as
@@
$\vec{e}_{\vec{k}\alpha,i}=\vec{p}_{\alpha}\exp\left(i\vec{k}\cdot\vec{r}_{0,i}\right),$ (53)
@@
$\xi_{L}=\sqrt{\frac{(K+G/3)\tau}{\gamma\rho}},\qquad\xi_{T}=\sqrt{\frac{G\tau}{\gamma\rho}}.$ (58)
These lengths can also be estimated in a different way. When we consider a transverse wave with the wavenumber $k$, the restoring force acting on this wave (per particle) in continuum mechanics is estimated as $Gk^{2}/\rho$, and its relaxation time is $\gamma\rho/(Gk^{2})$. If this relaxation time becomes longer than the persistence time tau, the quasi-static nature emerges. This gives an estimate of the characteristic wavenumber as $k_{T}\sim\sqrt{\gamma\rho/(G\tau)}$, which reproduces $\xi_{T}$ in Eq. (58). Likewise, for a longitudinal wave, we estimate the characteristic wavenumber as $k_{L}\sim\sqrt{\gamma\rho/[(K+G/3)\tau]}$, which reproduces $\xi_{L}$ in Eq. (58).

### v3

@@ I Introduction
When the active matter is in sufficiently dense states, it often exhibits glassy dynamics [6, 7, 8]. This phenomenon has received considerable attention in biological physics, especially in studies of the collective motion of confluent epithelial cells that regulate wound healing, cancer metastasis, and other processes [6, 9, 10, 11, 12, 13, 14, 15], the diffusion and viscoelasticity of proteins in bacterial cytoplasm [16, 17], and the dynamics and rheology of dense bacterial suspensions [18, 19]. Correspondingly, many numerical and theoretical models of dense active matter have been proposed and studied [20, 21, 22]. Among them, the simplest is presumably the self-propelled particles. Previous works have established that this model shows canonical glassy dynamics when the density is high enough or the active force is weak enough. In fact, the abrupt slowdown of the microscopic dynamics occers, the dynamic heterogeneity emerges, and the system eventually vitrifies into a glass state called the active glass. However, several non-canonical dynamical behaviors have also been observed, such as a change in the glass transition density, violation of the fluctuation-dissipation theorem, and the emergence of novel collective dynamics [20, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. Extensions of equilibrium theories of the glass transition have been proposed to understand these behaviors [33, 27, 30, 26, 34], but they are still the subject of intense debate.
@@
As mentioned above, previous studies have focused on the properties of dense active matter in their liquid state, while little is known about the properties of active glasses in their solid state. The present study aims to shed light on the structural fluctuations in active glasses. In passive glasses, particles are driven solely by thermal agitations, and their structural fluctuations are determined by their vibrational normal modes. Glasses have a large number of non-phonon normal modes, which differ significantly from crystalline solids where the normal modes are longitudinal and transverse phonons, and their vibrational density of states (vDOS) follows the Debye law $g(\omega)\propto\omega^{d-1}$ (where $d$ is the spatial dimension). As a result, there exist excess vibrational states over the Debye law, which is observed by a large peak in the reduced vDOS $g(\omega)/\omega^{d-1}$, referred to as the boson peak [35, 36, 37]. Additionally, at much lower frequencies, the vibrational normal modes become a mixture of phonon vibrations and quasi-localized vibrations [38, 39, 40, 41]. Within the harmonic description, structural fluctuations in passive glasses can be described in terms of the thermal excitations of these characteristic normal modes, all of which follow the equipartition law. Even in anharmonic dynamics at finite temperatures [42], the quasi-localized vibrations can be used to indicate structural rearrangements [43]. Therefore, to understand the structural fluctuations in active glasses, it is essential to examine how the active forces excite the characteristic normal modes in glasses.
@@
In the present work, we study active glasses by employing three different models: Active Ornstein Uhlenbeck Particles (AOUPs) [44], Active Brownian Particles (ABPs) [45], and Run and Tumble Particles (RTPs) [46]. To characterize structural fluctuations, we consider four observables: average excitation energy of the normal modes, mean-squared displacements (MSD), structure factor of longitudinal and transverse displacement fields, and non-Gaussian parameters (NGP). Firstly, we demonstrate that the former three observables can be described in terms of the two-time correlations of active forces and the normal modes of glasses. Next, we formulate the two-time correlations of the active forces to obtain general formulas of the three observables. Additionally, we prove that AOUPs, ABPs, and RTPs have the same two-time correlation when the active forces are properly reparameterized. We then apply the obtained formulas to calculate the observables for a typical configuration of active glasses. Our central observation is that active forces behave like random noise for low-frequency normal modes, while they behave like constant forces for high-frequency normal modes. This results in thermal excitations of low-frequency modes on one hand and quasi-static dynamics of high-frequency modes on the other, both of which determine nature of the structural fluctuations in active glasses. Finally, we study the NGP and find that fluctuations are mostly Gaussian, except for normal modes at the high-frequency edge. It should be noted that a previous study conducted by Henke et al. [47] investigated the velocity fields in confluent cells, both experimentally and theoretically. In their study, they modeled the cells using the ABPs, and demonstrated that the excitations of the normal modes in the model were able to accurately describe the velocity fields of cells as measured by experiments. The present work provides a more thorough analysis of the displacement fields, which complements and generalizes the work done by Henke et al. [47].
@@ II.2 Equation of motion
Because we assumed the mechanical stability of the configuration ${\bf\it r}_{0}$, the dynamical matrix $\mathcal{M}$ is positive semi-definite and has two translational zero modes. By removing these zero modes, we can focus on the fluctuation dynamics around ${\bf\it u}={\bf\it 0}$. In this study, we solve Eq. (5) to investigate the fluctuation dynamics in the steady state. We consider three different active forces, each corresponding to a different active glass model.
@@ II.2.1 Active Ornstein Uhlenbeck Particles
In the AOUPs [44], the active force follows
@@ II.2.2 Active Brownian Particles
In the ABPs [45], the particle tries to maintain a constant speed while their direction changes constantly. The active force acting on ABPs is described by
@@ II.2.3 Run and Tumble Particles
In the RTPs [46], the particle attempts to move at a constant speed, similar to the ABPs. However, their directions change abruptly and randomly. The force vector, ${\bf\it f}$, is described by Eq. (8). Meanwhile, each angle $\theta_{i}$ changes randomly in the range of $0\leq\theta_{i}<2\pi$ with a probability per unit time $P_{\text{\rm tumble}}=1-e^{-R_{\rm tumble}}$, where $R_{\rm tumble}$ is a constant rate. Hereafter, we refer to the active glass model with this active force as the RTP glass.
@@ II.4 Observables
Here, $\vec{u}_{\vec{k}}$ is the Fourier transform of the displacement field ${\bf\it u}=(\vec{u}_{1},\vec{u}_{2},\cdots,\vec{u}_{N})$, and when the displacement is sufficiently small, it is given by
$\vec{u}_{\vec{k}}=\sum_{i=1}^{N}\vec{u}_{i}\exp(-i\vec{k}\cdot\vec{r}_{0,i}),$ (18)
where $\vec{r}_{0,i}$ denotes the position of the particle $i$ in the inherent structure ${\bf\it r}_{0}$. The longitudinal and transverse parts are defined as $\vec{u}_{L,\vec{k}}=\left(\hat{\vec{k}}\cdot\vec{u}_{\vec{k}}\right)\hat{\vec{k}}$ and $\vec{u}_{T,\vec{k}}=\vec{u}_{\vec{k}}-\vec{u}_{L,\vec{k}}$, where $\hat{\vec{k}}=\vec{k}/k$ is the unit vector in the direction of $\vec{k}$, and $k=|\vec{k}|$. Using Eq. (12), we can express these correlation functions in terms of the eigenmodes as
@@
×$\times\left(\hat{\vec{k}}\cdot\vec{e}_{l,i}\right)\left(\hat{\vec{k}}\cdot\vec{e}^{\dagger}_{m,j}\right)\exp\left[i\left(\vec{k}\cdot\left(\vec{r}_{0,j}-\vec{r}_{0,i}\right)\right)\right],$
@@
×$\times\left[\vec{e}_{l,i}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}_{l,i}\right)\right]\left[\vec{e}^{\dagger}_{m,j}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}^{\dagger}_{m,j}\right)\right]$
@@
Finally, we quantify the non-Gaussianity of the fluctuation dynamics by introducing two NGPs. One NGP is defined for the active force as
@@ III.2 ABP glass
The time correlations on the right-hand side of Eq. (26) can be evaluated as follows. Since ${\bf\it\theta}(t)$ follows the Wiener process Eq.(9), the conditional probability $P({\bf\it\theta},t|{\bf\it\theta}^{\prime},t^{\prime})$, which expresses the probability that the angle vector is ${\bf\it\theta}$ at time $t$ given that it is ${\bf\it\theta}^{\prime}$ at time $t^{\prime}$, is expressed as [49]
@@
$=\gamma^{2}v_{0}^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\{e_{l,i}^{x}e_{m,j}^{x}+e_{l,i}^{y}e_{m,j}^{y}\}\frac{1}{2}e^{-D_{R}|t-t^{\prime}|}\delta_{ij},$
@@ III.3 RTP glass
The active force $f_{l}$ in the RTP glass shares the same expression Eq. (25) with the ABP glass. Therefore, the time correlations can be written as in Eq. (26). A difference between the RTPs and ABPs comes from the time evolution of ${\bf\it\theta}(t)$. In the RTP glass, ${\bf\it\theta}(t)$ follows the Poisson process, and the conditional probability is given by [49, 50]
@@ III.4 Comparison of three active forces
Moving forward, we will utilize the parameters $(D,\tau)$ to identify the three active forces. Specifically, $(v_{0}=\sqrt{2D/\tau\gamma},D_{R}=1/\tau)$ is used to describe the ABP glass, while $(v_{0}=\sqrt{2D/\tau\gamma},R_{\rm tumble}=1/\tau)$ is used to describe the RTP glass. This parameterization will enable us to compare the three active glasses on an equal footing of $(D,\tau)$.
@@ IV Numerical model
In addition, we conduct numerical simulations by solving the linearized equation of motion, Eq. (5), with three different active forces, all of which begin from the same mechanically stable configuration ${\bf\it r}_{0}$. We adopted the Predictor-Corrector Method for the numerical integration. The active forces were computed using different methods depending on the type of system. For the AOUP glass, we use the Gillespie method [53], while for the ABP glass, we use the Euler-Maruyama method. For the RTP glass, we select the update event of the active forces to occur with a probability $P_{\text{tumble}}=1-e^{-R_{\text{tumble}}\Delta t}$ at each time step $\Delta t$, and then the direction of the active forces was chosen according to the uniform distribution. In all three cases, we perform simulations for a sufficiently long period to achieve stationary states before running the production runs.
@@ V.1 Time correlation function of $c_{l}$
This outcome is identical to the one for thermal glasses [54], where $D$ denotes temperature and $\gamma\omega_{l}^{2}$ symbolizes the stiffness of the mode $l$. It demonstrates that under the condition $\tau\ll\omega_{l}^{-2}$ which we term the “thermal condition” from now on, the active forces function like a random white noise for the mode $l$, and the dynamics along it are equivalent to those in thermal systems.
@@
It is important to note that the typical amplitude of $c_{l}$ is heavily influenced by these conditions, as discussed in detail later. In the thermal condition, we observe that $\langle c_{l}^{2}\rangle\propto\omega_{l}^{-2}$, which is consistent with the equipartition law. However, in the quasi-static condition, $\langle c_{l}^{2}\rangle\propto\omega_{l}^{-4}$, meaning that displacements along the high-frequency modes are significantly suppressed.
@@ V.4 Spatial correlation of displacements
We now discuss the spatial correlations of the longitudinal and transverse displacement fields, denoted as $S_{L}(k)$ and $S_{T}(k)$, respectively. These correlation functions are calculated in our numerical simulations using Eq. (17). In Figs. 4(a) and 5(a), we present $S_{L}(k)$ and $S_{T}(k)$ for three different active glasses with varying persistent times $\tau$. When $\tau=10^{-8}$, $S_{L}(k)$ and $S_{T}(k)$ exhibit a plateau in a wide range of $k$. This is consistent with the known results for thermal glasses [48]. However, as $\tau$ is increased, both $S_{L}(k)$ and $S_{T}(k)$ become smaller at large $k$, indicating that short-wavelength excitations are strongly suppressed. This indicates that the fluctuations of active glasses exhibit apparent collective motions as the long-wavelength excitations are emphasized [47]. These apparent collective motions only appear when the persistence time $\tau$ is sufficiently large. Notably, the above observations for all three active glasses are identical.
@@
×$\times\left(\hat{\vec{k}}\cdot\vec{e}_{l,i}\right)\left(\hat{\vec{k}}\cdot\vec{e}^{\dagger}_{l,j}\right)\exp\left[i\left(\vec{k}\cdot\left(\vec{r}_{0,j}-\vec{r}_{0,i}\right)\right)\right],$
@@
×$\times\left[\vec{e}_{l,i}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}_{l,i}\right)\right]\left[\vec{e}^{{\dagger}}_{l,j}-\hat{\vec{k}}\left(\hat{\vec{k}}\cdot\vec{e}^{{\dagger}}_{l,j}\right)\right]$
@@
Here, we focus on the low-wavenumber regime in Eq. (50). We can use conventional elasticity theory to approximate the eigenmodes in this regime [55]. Specifically, in the two-dimensional continuum solid, we can approximate the dynamical matrix Eq. (3) as the operator ${\frac{1}{\gamma\rho}\left(K\nabla\nabla+G\nabla^{2}\right)}$, which acts on the continuum displacement fields. Here, $K$ and $G$ are the bulk and shear modulus, respectively, and $\rho=N/V$ is the number density. We can solve the eigenvalue problem of this operator and obtain the eigenmodes specified by $l=\vec{k}\alpha$ where $\vec{k}$ is the wavevector and $\alpha$ is the polarization with $\alpha=L$ and $T$ denoting the longitudinal and the transverse modes, respectively. The eigenfrequency can be expressed as
@@
$\vec{e}_{\vec{k}\alpha,i}=\frac{\vec{p}_{\alpha}}{\sqrt{N}}\exp\left(i\vec{k}\cdot\vec{r}_{0,i}\right),$ (53)
@@
$\xi_{L}=\sqrt{\frac{(K+G)\tau}{\gamma\rho}},\qquad\xi_{T}=\sqrt{\frac{G\tau}{\gamma\rho}}.$ (58)
These lengths can also be estimated in a different way. When we consider a transverse wave with the wavenumber $k$, the restoring force acting on this wave (per particle) in continuum mechanics is estimated as $Gk^{2}/\rho$, and its relaxation time is $\gamma\rho/(Gk^{2})$. If this relaxation time becomes longer than the persistence time $\tau$, the quasi-static nature emerges. This gives an estimate of the characteristic wavenumber as $k_{T}\sim\sqrt{\gamma\rho/(G\tau)}$, which reproduces $\xi_{T}$ in Eq. (58). Likewise, for a longitudinal wave, we estimate the characteristic wavenumber as $k_{L}\sim\sqrt{\gamma\rho/[(K+G)\tau]}$, which reproduces $\xi_{L}$ in Eq. (58).